OpenStudy (anonymous):

At what altitude will a person weighs 150 lb become 75 lb.. help plss

OpenStudy (anonymous):

Well, the weight of an object is inversely proportional to the distance sqaured. According to Newton's Law of Graviation, F = (G * m1 * m2) / (r^2), in order to cut the force in half, the distance has to increase by the square root of 2. Earth's radius is 3963 miles, so 1.41 times that is 5604 miles. But that's the distance away from the core. To get the altitude, subtract earth's radius. You get an answer of 1641 miles above Earth's surface. Hope this helped!

OpenStudy (anonymous):

he answer should be 2650 km.. i dont know the soution

OpenStudy (anonymous):

This would be easier if the units were in consistent systems. Actually, the world would be better if we just ditched imperial altogether for metric. </rant> The units for mass in Imperial is slugs, so the person's mass on earth is going to be their weight divided by the acceleration due to gravity, i.e., \[m_e=\frac{150lbs}{32\frac{ft}{s^2}}\] Of course your final answer is supposed to be in km, so I'd start by finding their mass in kg and converting the other units as well before you get started. It'll be a *lot* easier.